A bank offers a savings account that pays interest quarterly with an effective annual rate (EAR) of 14%. What is the applicable quarterly periodic interest earned in this account?

a.

3.50%

b.

6.77%

c.

13.87%

d.

3.33%

e.

4.46%
Answer #1

EAR = (1-6)*-1 i = stated annual interest rate n = number of compounding periods

0.14 = (1 + \frac{i}{4})^4 - 1

1.14 = (1 + \frac{i}{4})^4

1.0333 = 1 + (i/4)

(i/4) = 3.33% --> APplicable quarterly interest rate (Option d)